OpenStudy (anonymous):

The length of a rectangle is 7 m less than triple the width. If w represents the width, write an expression for the length, l. A. B. l = 3w – 7 C. D. l = 3w + 7 I am not good with problems like this

OpenStudy (anonymous):

A. l=1/3w-7 C. l=1/3w+7

OpenStudy (anonymous):

it is B

OpenStudy (anonymous):

B I=3w-7

OpenStudy (anonymous):

thank you